# Stacked Intelligent Metasurface-Enhanced Wideband Multiuser MIMO OFDM-IM Communications

### v1

@@ Abstract
Leveraging the multilayer realization of programmable metasurfaces, stacked intelligent metasurfaces (SIM) enable fine-grained wave-domain control. However, their wideband deployment is impeded by two structural factors: (i) a single, quasi-static SIM phase tensor must adapt to all subcarriers, and (ii) multiuser scheduling changes the subcarrier activation pattern frame by frame, requiring rapid reconfiguration. To address both challenges, we develop a SIM-enhanced wideband multiuser transceiver built on orthogonal frequency-division multiplexing with index modulation (OFDM–IM). The sparse activation of OFDM-IM confines high-fidelity equalization to the active tones, effectively widening the usable bandwidth. To make the design reliability-aware, we directly target the worst-link bit-error rate (BER) and adopt a max-min per-tone signal-to-interference-plus-noise ratio (SINR) as a principled surrogate, turning the reliability optimization tractable. For frame-rate inference and interpretability, we propose an unfolded projected-gradient-descent network (UPGD-Net) that double-unrolls across the SIM’s layers and algorithmic iterations: each cell computes the analytic gradient from the cascaded precoder with a learnable per-iteration step size. Simulations on wideband multiuser downlinks show fast, monotone convergence, an evident layer-depth sweet spot, and consistent gains in worst-link BER and sum rate. By combining structural sparsity with a BER-driven, deep-unfolded optimization backbone, the proposed framework directly addresses the key wideband deficiencies of SIM.
@@ I Introduction
Recent studies have demonstrated SIM’s versatility across a range of applications. For instance, [17, 18] illustrated SIM’s potential for highly controllable signal transmission and reception in holographic MIMO systems. SIM-based beamforming techniques were explored in [19, 20, 21] to demonstrate their superiority over conventional MIMO in energy efficiency. Multiuser communication performance enhancements by SIMs were reported in [13, 22, 23, 24, 25, 26], emphasizing their adaptability and effectiveness. SIM has also been applied to direction-of-arrival (DoA) estimation [27], satellite communication [28], communication security [29], and semantic communication [30]. In both cellular and cell-free networks, SIM has shown promise in mitigating inter-user interference (IUI) and enhancing resource allocation [31, 32, 33]. Moreover, acting like an artificial neural network (ANN), SIM packs multiple layers of metasurfaces for advanced computation and signal processing tasks, with the configured transmission coefficients serving as trainable network parameters. Therefore, SIM can be employed to support many beyond-communication functionalities, such as wireless sensing and localization [34, 35, 36], while recent efforts have investigated reinforcement learning and meta-learning frameworks to dynamically configure SIM for various scenarios [37, 38]. In summary, most previous works on SIM focus on narrowband systems with frequency-flat channels [39, 40], where analog signal processing can be easily tackled with high energy efficiency in the wave domain.
@@
OFDM with index modulation (OFDM–IM) provides the required structural complement for wideband SIM applications from the coding and modulation side. In each OFDM-IM subblock, only a subset of subcarriers is activated, and their positions, which are determined by the incoming index bits, serve as an additional information-bearing dimension [43, 44, 45]. The induced spectral sparsity (i) lowers peak-to-average-power ratio (PAPR) by reducing the probability of large-symbol superposition in time [46], and (ii) requires high-quality equalization only on the active tones, thereby effectively widening the bandwidth over which a given SIM can realize fully–analog beamforming. Specifically, the reduced number of active subcarriers per user significantly alleviates the burden of jointly optimizing high-dimensional SIM phase-shift matrices across a wideband spectrum. Prior work [1] quantified this trade-off of activation pattern selections in a SIM-enhanced multiuser system, which formulated the channel fitting optimization problem with subcarrier allocation to relax the interference-mitigation constraints. In addition, RIS-aided OFDM-IM designs were proposed in [47, 48] to leverage the subcarrier allocation to transmit extra information efficiently, where phase shifts were configured by an exhaustive search-based optimization algorithm. However, no SIM-enhanced multiuser OFDM–IM transceiver benchmarked against a bit error rate (BER) metric has been established. Bridging this gap advances the SIM designs from narrowband/rate-centric formulations to a wideband, reliability-aware, fully-analog architecture.
@@
To fill the above research gaps, we develop a SIM-enhanced multiuser OFDM-IM transceiver and a physics-guided deep-unfolding solver, tailored to wideband communications. A multilayer SIM performs wave-domain zero-forcing (ZF) to mitigate IUI, while OFDM-IM supplies frequency-domain sparsity that lowers PAPR and relaxes the equalization burden. By grouping nearest-neighbor error events into distance classes and introducing a single distance normalizer, we establish a compact bridge from the worst-link BER to a distance-weighted per-tone signal-to-interference-plus-noise-ratio (SINR). This converts BER minimization exactly into a max-min SINR problem that collapses to a single second-order cone programming problem, enabling joint phase-power-index design. The projected-gradient-descent (PGD) for this program is then unfolded into a double-unrolled network (iterations $\times$ layers), preserving full interpretability. Each cell executes a closed-form PGD update with unit-modulus projection, while only the per-iteration step sizes are learned offline to avoid overfitting. And the training database can cover hundreds of channel/activation patterns.
@@
  • We derive an analytical BER-to-SINR bridge under SIM-enhanced beamforming and convert BER minimization into a distance-weighted max-min SINR problem. The problem reformulation turns a non-convex BER objective into an implementation-ready tool for joint phase, power, and index allocation, closing the gap between SINR heuristics and true error-rate optimization.
  • We design a double-unrolled PGD network (UPGD-Net) aligned with SIM physics: unrolling across the SIM’s layers and iterations, using analytic worst-link SINR gradients, unit-modulus projection, and a small set of learnable step sizes per iteration. The solver preserves hard constraints and monotone descent, offers interpretability, and enables per-frame inference once trained.
@@ II-C Wideband Channel and Received Signal Model
We consider a wideband multipath channel model, incorporating multiple scatterers that result in frequency-selective fading effects. The tapped delay line channel accounts for $P_{k}$ resolvable paths between the SIM at the BS and the $k$-th UE, where the complex gain and delay associated with the $p$-th path are denoted as $g_{p}$ and $\tau_{p}$, respectively. The frequency-domain channel vector on subcarrier $i$ is
$\mathbf{h}_{k}(i)=\sum_{p=0}^{P_{k}}g_{p}(f_{i})e^{-j2\pi f_{i}\tau_{p}}{\bm{\alpha}_{p}(i)}^{H}\in\mathbb{C}^{1\times M},$ (6)
@@ III-A Single-Subblock BER Derivation
Let $\mathbf{x}_{k,\ell}$ denote the transmitted symbol in the frequency domain and $\widehat{\mathbf{x}}_{k,\ell}$ denote the signal detected erroneously by the RX in valid indices and constellation symbols. For a fixed channel under the TDD protocol, the conditional pairwise error probability (CPEP) in [56] collapses to the deterministic pairwise error probability (PEP) form between two OFDM-IM blocks $\mathbf{x}_{k,\ell}$ and $\widehat{\mathbf{x}}_{k,\ell}$ of UE $k$ in the $\ell$-th subblock
@@ III-B Network-Level Worst‑Link BER and Distance‑Weighted SINR
Denote $P_{b,k}$ the BEP over the entire frame. Because the $L$ subblocks are statistically independent and carry the same number of information bits, we arrive at
@@
We adopt the worst-link BER as the quality of service (QoS) target across the frame,
@@
Define the largest distance $\beta_{\max}=\max_{c\in\mathcal{C}}\beta_{c}$ and the normalized SINR $\eta_{c}=\frac{\gamma_{\min}^{(c)}}{\beta_{\max}}$. Multiply and divide the product $\beta_{c}\,\gamma_{\min}^{(c)}$ by $\beta_{\max}$:
$\beta_{c}\,\gamma_{\min}^{(c)}\,=\,\beta_{c}\,\frac{\beta_{\max}}{\beta_{\max}}\,\gamma_{\min}^{(c)}\,\geq\,{\beta_{\max}\,{\frac{\gamma_{\min}^{(c)}}{\beta_{\max}}}},$ (22)
Take the minimum over the four classes $c\in\mathcal{C}$:
$\eta\;=\;\min_{c\in\mathcal{C}}\eta_{c}\;=\;\min_{c\in\mathcal{C}}\frac{\gamma_{\min}^{(c)}}{\beta_{\max}}\;=\;\frac{1}{\beta_{\max}}\min_{n\in\mathcal{N}}\gamma_{k,\ell,n}.$ (23)
Every error class $c$ contributes an error exponent $E_{k,\ell,c}$ to the union bound. With (22) and (23), replacing $\Sigma_{k,\ell,c}$ by its worst–link lower bound and upper‑bounding the coefficient $\beta_{c}\leq\beta_{\max}$ gives
$E_{k,c}\;\leq\;\exp\!\Bigl(-\tfrac{\beta_{\max}\,\eta_{c}}{4\sin^{2}\zeta}\,\Upsilon_{k,c}\Bigr)\;\leq\;\exp\!\Bigl(-\tfrac{\beta_{\max}\,\eta}{4\sin^{2}\zeta}\,\Upsilon_{k,c}\Bigr).$ (24)
Since $E_{k,\ell,c}$ is decreasing monotonically in $\eta$, the worst-link BEP is also monotonically decreasing in $\eta$.
Minimizing the worst‑link BEP is therefore equivalent to maximizing $\eta$
$\min_{\theta_{m}^{l}}\;P_{b}^{\max}\quad\Longleftrightarrow\quad\max_{\theta_{m}^{l}}\;\eta=\frac{1}{\beta_{\max}}\min_{k,\ell,n}\gamma_{k,\ell,n}.$ (25)
@@
Without the scaling in (24), we would obtain three various class-specific squared Euclidean distances $\eta_{\textbf{C}_{1}},\eta_{\textbf{C}_{2}},\eta_{\textbf{C}_{3}}$ and hence three separate inequality sets. It increases the number of SINR constraints by a factor of three while delivering the same numerical BER (the largest distance class dominates) to increase the computational complexity. The normalization factor $\beta_{\max}$ compresses all of them into one constraint, keeps the optimization dimension minimal, and aligns the effective distance squared across classes.
@@
All class-dependent squared Euclidean distances $\beta_{c}$, pair multiplicity $N_{k,\ell,c}$, and bit error weight $w_{k,\ell,c}$ are determined solely by the modulation alphabet $\mathbf{S}_{k,\ell}$ and the OFDM‑IM activation rule $(N,V)$, which do not depend on channel, precoder, or power allocation. Once fixed offline, these parameters act as static scaling factors within the SINR constraints and therefore do not enlarge the online optimization dimension.
@@ III-C Final Optimization Problem
$\textstyle\sum_{k=1}^{K}Z_{k,\ell,n}\,\geq\,1,~\forall\ell\in\mathcal{L}_{b},\;\forall n\in\mathcal{N},$ (27d)
$\mathbf{Z}\in\{0,1\}^{K\times N_{c}},$ (27e)
The optimization problem formulated in Section III aims to maximize the worst-link SINR across all UEs and active subcarriers by tuning the phase shifts of a fully-analog SIM-enhanced precoder. However, it is challenging to solve the max-min problem due to the nonlinear and highly coupled structure of the SINR expression w.r.t. $\theta_{m}^{l}$}. Furthermore, in OFDM-IM, each transmitted bitstream activates a distinct subset of subcarriers according to the index bits, which corresponds to a unique subcarrier activation matrix $\mathbf{Z}$. In traditional solvers, every change in the subcarrier activation matrix $\mathbf{Z}$ requires solving the full max-min SINR optimization problem in (27) from scratch. This is computationally expensive and impractical for real-time OFDM-IM transmission where $\mathbf{Z}$ may be frequently re-optimized due to UE scheduling or adaptive resource allocation. To address this challenge, we propose a DUN that embeds the structure of the PGD method into a trainable deep neural architecture in Section IV to obtain a SIM phase solution for problem $\mathcal{P}_{2}$.

### v2

@@ Abstract
Stacked intelligent metasurfaces (SIM) enable fine-grained wave-domain signal processing, but their wideband deployment is impeded by two structural factors: (i) a single, quasi-static SIM phase tensor must adapt to all subcarriers, and (ii) multiuser scheduling changes the subcarrier activation pattern frame by frame, requiring rapid reconfiguration. To address these, we propose a SIM-enhanced wideband multiuser transceiver built on orthogonal frequency-division multiplexing with index modulation (OFDM–IM). The sparse activation of OFDM-IM confines high-fidelity equalization to the active tones, effectively widening the usable bandwidth. To make the design reliability-aware, we directly target the worst-link bit-error rate (BER) and adopt a max-min per-tone signal-to-interference-plus-noise ratio (SINR) as a principled surrogate, turning the reliability optimization tractable. For frame-rate inference and interpretability, we propose an unfolding projected-gradient-descent network (UPGD-Net) that unrolls across the SIM’s layers and algorithmic iterations with a learnable per-iteration step size. Simulations demonstrate that the proposed framework achieves fast convergence and significant BER gains over fully-digital baselines. Notably, the design exhibits superior robustness against errors and outperforms large-aperture hybrid precoding benchmarks in both sum rate and energy efficiency. By combining structural sparsity with a BER-driven, deep-unfolded optimization backbone, the proposed framework effectively resolves the key wideband deficiencies of SIM.
@@ I Introduction
Recent studies have demonstrated SIM’s versatility across a range of applications. For instance, [17, 18] illustrated SIM’s potential for highly controllable signal transmission and reception in holographic MIMO systems. SIM-based beamforming techniques were explored in [19, 20, 21] to demonstrate their superiority over conventional MIMO in energy efficiency. Multiuser communication performance enhancements by SIMs were reported in [13, 22, 23, 24, 25, 26], emphasizing their adaptability and effectiveness. SIM has also been applied to direction-of-arrival (DoA) estimation [27], satellite communication [28], communication security [29], and semantic communication [30]. In both cellular and cell-free networks, SIM has shown promise in mitigating inter-user interference (IUI) and enhancing resource allocation [31, 32, 33]. Moreover, acting like an artificial neural network (ANN), SIM packs multiple layers of metasurfaces for advanced computation and signal processing tasks, with the configured transmission coefficients serving as trainable network parameters. Therefore, SIM can be employed to support many beyond-communication functionalities, such as wireless sensing and localization [34, 35, 36], while recent efforts have investigated reinforcement learning and multi-agent frameworks to dynamically configure SIM for various scenarios [37, 38]. In summary, most previous works on SIM focus on narrowband systems with frequency-flat channels [39, 40], where analog signal processing can be easily tackled with high energy efficiency in the wave domain.
@@
OFDM with index modulation (OFDM–IM) provides the required structural complement for wideband SIM applications from the coding and modulation side. In each OFDM-IM subblock, only a subset of subcarriers is activated, and their positions, which are determined by the incoming index bits, serve as an additional information-bearing dimension [43, 44, 45]. The induced spectral sparsity (i) lowers peak-to-average-power ratio (PAPR) by reducing the probability of large-symbol superposition in time [46], and (ii) requires high-quality equalization only on the active tones, thereby effectively widening the bandwidth over which a given SIM can realize fully–analog beamforming. Specifically, since the subcarrier activation set is a subset of the full spectrum, optimizing high-dimensional SIM phase-shift matrices in OFDM-IM is less constrained than in full-carrier systems. This effectively relaxes the wideband beamforming requirements, allowing the frequency-static SIM phases to achieve higher equalized gain on the targeted active tones. Prior work [1] quantified this trade-off of activation pattern selections in a SIM-enhanced multiuser system, which formulated the channel fitting optimization problem with subcarrier allocation to relax the interference-mitigation constraints. In addition, RIS-aided OFDM-IM designs were proposed in [47, 48] to leverage the subcarrier allocation to transmit extra information efficiently, where phase shifts were configured by an exhaustive search-based optimization algorithm. However, no SIM-enhanced multiuser OFDM–IM transceiver benchmarked against a bit error rate (BER) metric has been established. Bridging this gap advances the SIM designs from narrowband/rate-centric formulations to a wideband, reliability-aware, fully-analog architecture.
@@
To fill the above research gaps, we develop a SIM-enhanced multiuser OFDM-IM transceiver and a physics-guided deep-unfolding solver, tailored to wideband communications. A multilayer SIM performs wave-domain zero-forcing (ZF) to mitigate IUI, while OFDM-IM supplies frequency-domain sparsity that lowers PAPR and relaxes the equalization burden. By grouping nearest-neighbor error events into distance classes and introducing a single distance normalizer, we establish a compact bridge from the worst-link BER to a distance-weighted per-tone signal-to-interference-plus-noise-ratio (SINR). This converts BER minimization into a principled surrogate, formulated as a constrained max-min SINR optimization problem. The projected-gradient-descent (PGD) for this program is then unfolded into a coupled physical-algorithmic network, preserving full interpretability. Each cell executes a closed-form PGD update with unit-modulus projection, while only the per-iteration step sizes are learned offline to avoid overfitting. And the training database can cover hundreds of channel/activation patterns.
@@
  • We derive an analytical BER-to-SINR bridge under SIM-enhanced beamforming and convert BER minimization into a distance-weighted max-min SINR problem. The problem reformulation turns a non-convex BER objective into a tractable surrogate for index-aware phase optimization, closing the gap between SINR heuristics and true error-rate optimization.
  • We design a coupled physical-algorithmic unfolding PGD network (UPGD-Net) aligned with SIM physics: unrolling across the SIM’s layers and iterations, using analytic worst-link SINR gradients, unit-modulus projection, and a small set of learnable step sizes per iteration. The solver preserves hard constraints and monotone descent, offers interpretability, and enables per-frame inference once trained.
@@ II-A SIM–Enhanced OFDM-IM Transceiver Architecture
To establish a fundamental performance benchmark for the proposed SIM-enhanced transceiver, we adopt the assumption of continuous phase shifts with a unit-modulus amplitude response for each meta-atom [56]. This modeling choice not only facilitates the derivation of a theoretical upper bound but also ensures the differentiability required for optimization. While practical implementation relies on discrete phases, recent studies [25] have validated the inherent robustness of SIM architectures to quantization errors. Specifically, a coarse 2-bit phase resolution incurs only a marginal sum rate penalty (<0.5<0.5 bps) compared to the continuous profile, with the performance gap becoming negligible at higher resolutions ($\geq 3$ bits). Consequently, the continuous model serves as a rigorous baseline for the wave-domain beamforming principles explored in this study. Detailed investigations into other hardware non-idealities, such as adjustable magnitudes [5] and coupled phase-amplitude tuning mechanisms [4], are reserved for future work.
Remark 2.
@@ II-C Wideband Channel and Received Signal Model
We consider a wideband multipath channel model, incorporating multiple scatterers that result in frequency-selective fading effects. The tapped delay line channel accounts for $P_{k}$ resolvable paths between the SIM at the BS and the $k$-th UE. Let $p=0$ denote the line-of-sight (LoS) path and $p=1,\ldots,P_{k}$ denote the non-LoS (NLoS) paths. The complex gain and delay associated with the $p$-th path are denoted as $g_{p}$ and $\tau_{p}$, respectively. The frequency-domain channel vector on subcarrier $i$ is
${{\mathbf{h}_{k}(i)=\sum_{p=0}^{P_{k}}g_{p}e^{-j2\pi f_{i}\tau_{p}}{\bm{\alpha}_{p}(i)}^{H}\in\mathbb{C}^{1\times M},}}$ (6)
@@
Remark 5.
The validity of the forward-propagation model is verified by the specific hardware architecture adopted in this work, consistent with the transmissive SIM prototypes reported in [11]. By utilizing meta-atoms optimized for unidirectional transmission, the system inherently suppresses reverse coupling between layers and self-oscillation, ensuring that electromagnetic energy propagates primarily in the forward direction. Structurally, the metasurface stack is enclosed within a shielded metallic chassis lined with absorptive materials to mitigate edge/side diffraction and lateral leakage. Furthermore, the cumulative energy of higher-order reflections between layers attenuates exponentially, and an additional anti-reflection coating is integrated at the back of the stack to eliminate residual backward scattering. As such, we do not consider backward and anomalous waves in our theoretical model. Nonetheless, future designs could intentionally leverage controlled inter-layer feedback to realize wave-domain recursive networks.
Remark 6.
We acknowledge that the analytical transmission models in (1), (2) and (6) may exhibit discrepancies when compared to practical hardware implementations. These deviations arise not only from inevitable fabrication tolerances but also from simplified physical assumptions, specifically the omission of resistive/dielectric losses and intra-layer mutual coupling effects. To bridge this gap, an in-situ calibration procedure is essential before the SIM’s practical deployment. By transmitting pilot sequences and measuring the received response for the end-to-end learning, the effective transmission coefficients can be iteratively refined using error back-propagation algorithms [57]. This data-driven update naturally compensates for the aforementioned modeling mismatches and hardware imperfections. Given that this work focuses on the fundamental precoding framework in MIMO OFDM-IM systems, the detailed implementation of such calibration schemes is designated for future work.
@@ III-A Single-Subblock BER Derivation
Let $\mathbf{x}_{k,\ell}$ denote the transmitted symbol in the frequency domain and $\widehat{\mathbf{x}}_{k,\ell}$ denote the signal detected erroneously by the RX in valid indices and constellation symbols. Here $x_{k,\ell,n},\widehat{x}_{k,\ell,n}\in\mathbf{S}_{k,\ell}\cup\{0\}$, where $\mathbf{S}_{k,\ell}$ denotes the modulation alphabet and 0 models an inactive tone in OFDM-IM. For a fixed channel under the TDD protocol, the conditional pairwise error probability (CPEP) in [58] collapses to the deterministic pairwise error probability (PEP) form between two OFDM-IM blocks $\mathbf{x}_{k,\ell}$ and $\widehat{\mathbf{x}}_{k,\ell}$ of UE $k$ in the $\ell$-th subblock
@@ III-B Network-Level Worst‑Link BER and Distance‑Weighted SINR
Let $P_{b,k}$ denote the BEP over the entire frame. Because the $L$ subblocks are statistically independent and carry the same number of information bits, we arrive at
@@
We adopt the worst-case BEP among all users as the optimization objective to ensure fairness:
@@
Define the minimum distance parameter $\beta_{\min}=\min_{c\in\mathcal{C}}\beta_{c}$ and $\eta_{c}=\gamma_{\min}^{(c)}/\beta_{\min}$. By scaling the effective signal power term $\beta_{c}\,\gamma_{\min}^{(c)}$ using the lower bound $\beta_{c}\geq\beta_{\min}$, we obtain:
$\beta_{c}\,\gamma_{\min}^{(c)}=\beta_{c}\,\frac{\beta_{\min}}{\beta_{\min}}\,\gamma_{\min}^{(c)}\ \geq\ \beta_{\min}^{2}\,\frac{\gamma_{\min}^{(c)}}{\beta_{\min}}.$ (22)
Taking the minimum over the three classes $c\in\mathcal{C}$ gives
$\eta=\min_{c\in\mathcal{C}}\eta_{c}=\min_{c\in\mathcal{C}}\frac{\gamma_{\min}^{(c)}}{\beta_{\min}}=\frac{1}{\beta_{\min}}\min_{n\in\mathcal{N}}\gamma_{k,\ell,n}.$ (23)
Each class $c$ contributes an exponent $E_{k,\ell,c}$ to the union bound. Lower-bounding $\Sigma_{k,\ell,c}$ by its worst–link term and using $\beta_{c}\geq\beta_{\min}$ (replacing $\beta_{c}$ by $\beta_{\min}$) yield
$E_{k,\ell,c}\ \leq\ \exp\!\Bigl(-\tfrac{\beta_{\min}^{2}\,\eta_{c}}{4\sin^{2}\zeta}\,\Upsilon_{k,\ell,c}\Bigr)\ \leq\ \exp\!\Bigl(-\tfrac{\beta_{\min}^{2}\,\eta}{4\sin^{2}\zeta}\,\Upsilon_{k,\ell,c}\Bigr).$ (24)
Since each $E_{k,\ell,c}$ is a decreasing function of $\eta$, and summation over $c$ and integration over $\zeta$ preserve monotonicity (all weights are positive), the worst-link union-bound BER is also monotonically decreasing with respect to $\eta$.
Therefore, the original BEP minimization problem is cast into the following surrogate optimization problem:
$\min_{\theta_{m}^{l}}\;P_{b}^{\max}\quad\xrightarrow{\text{Surrogate}}\quad\max_{\theta_{m}^{l}}\;\eta=\frac{1}{\beta_{\min}}\min_{k,\ell,n}\gamma_{k,\ell,n}.$ (25)
@@
Without the normalization strategy in (24), the optimization would require three distinct inequality constraints corresponding to the class-specific distances $\eta_{\textbf{C}_{1}},\eta_{\textbf{C}_{2}},\eta_{\textbf{C}_{3}}$. This would triple the constraint volume of the SINR sub-problem while yielding the same numerical BER performance (as the process is dominated by the minimum distance). The introduction of $\beta_{\min}$ compresses these requirements into a single scalar constraint, minimizing the optimization dimensionality and aligning the effective Euclidean distance across different error types.
@@
The class-dependent parameters, including the squared Euclidean distances $\beta_{c}$, pair multiplicity $N_{k,\ell,c}$, and bit error weights $w_{k,\ell,c}$, are intrinsic properties determined solely by the modulation alphabet $\mathbf{S}_{k,\ell}$ and the OFDM‑IM activation rule $(N,V)$. These parameters are independent of the channel realization, precoder design, or power allocation. Consequently, they are pre-calculated offline and act as static scaling factors, introducing no additional complexity to the online optimization process.
@@ III-C Final Optimization Problem
$\mathbf{Z}\in\{0,1\}^{K\times N_{c}},$ (27d)
The optimization problem formulated in Section III aims to maximize the worst-link SINR across all UEs and active subcarriers by tuning the phase shifts of a fully-analog SIM-enhanced precoder. $\mathbf{Z}$ is a fixed input parameter determined by the data source, and the optimization is performed over $\theta_{m}^{l}$} conditioned on $\mathbf{Z}$. However, it is challenging to solve the max-min problem due to the nonlinear and highly coupled structure of the SINR expression w.r.t. $\theta_{m}^{l}$}. Furthermore, in OFDM-IM, each transmitted bitstream activates a distinct subset of subcarriers according to the index bits, which corresponds to a unique subcarrier activation matrix $\mathbf{Z}$. In traditional solvers, every change in the subcarrier activation matrix $\mathbf{Z}$ requires solving the full max-min SINR optimization problem in (27) from scratch. This is computationally expensive and impractical for real-time OFDM-IM transmission where $\mathbf{Z}$ may be frequently re-optimized due to UE scheduling or adaptive resource allocation. To address this challenge, we propose a DUN that embeds the structure of the PGD method into a trainable deep neural architecture in Section IV to obtain a SIM phase solution for problem $\mathcal{P}_{2}$.